OpenStudy (anonymous):

r^2-22r+21=0 How am I suppose to start this off? Kinda having a brain lag at this time of night

OpenStudy (anonymous):

you have several ways to solve it 1) grouping method 2)by completing square 3) by using direct formula which one you like

OpenStudy (anonymous):

grouping

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

let me try that

OpenStudy (anonymous):

ok try it

OpenStudy (chmvijay):

by grouping it may not be possible i think

OpenStudy (anonymous):

it possible

OpenStudy (anonymous):

im trying

OpenStudy (anonymous):

yeah thats not working out

OpenStudy (anonymous):

can i do it for you??

OpenStudy (anonymous):

if I group i get

OpenStudy (chmvijay):

yaaa please do it

OpenStudy (anonymous):

ok see

OpenStudy (anonymous):

r^2-1r-21r+21 Is this correct?

OpenStudy (anonymous):

r^2-22r+21=0 r^2-21r-r+21=0 r(r-21)-1(r-21)=0 (r-1)(r-21)=0 => r = 1 or r = 21

OpenStudy (chmvijay):

yaaa its possible

OpenStudy (chmvijay):

great job

OpenStudy (anonymous):

Damn good job!

OpenStudy (anonymous):

So therefore the sol set is 1 & 21?

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

great

OpenStudy (anonymous):

What it an easier way of figuring this equation out without using the quadratic formula? z^2+5z=-6

OpenStudy (anonymous):

I got it!

OpenStudy (anonymous):

oooooooooooooooooooook

OpenStudy (anonymous):

z=-3,-2

OpenStudy (anonymous):

yessssss

OpenStudy (anonymous):

awesome!

OpenStudy (anonymous):

you are good at maths. good try

OpenStudy (anonymous):

Thanks Im trying

OpenStudy (anonymous):

you are welcome

OpenStudy (anonymous):

s^2=9 I got 3,-3

OpenStudy (anonymous):

yes thats right

OpenStudy (anonymous):

good deal

OpenStudy (anonymous):

How should I start this one off? 3b^2=9b

OpenStudy (anonymous):

ok divide both sides by b first

OpenStudy (anonymous):

k

OpenStudy (anonymous):

b=3?

OpenStudy (anonymous):

yap

OpenStudy (anonymous):

But its asking for the solutions sets?

OpenStudy (anonymous):

ok ok bring that 9b to LHS what you get

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

9b-3b^2=0??

OpenStudy (anonymous):

its 3b^2-9b = 0 take 3b common what you get???

OpenStudy (anonymous):

take 3b^2 huh?

OpenStudy (anonymous):

3b(b-3)

OpenStudy (anonymous):

yessss

OpenStudy (anonymous):

3b=0 and b-3=0

OpenStudy (anonymous):

thats right so what you get for b???

OpenStudy (anonymous):

0,3

OpenStudy (anonymous):

thats right

OpenStudy (anonymous):

sweet! Thanks

OpenStudy (anonymous):

ok for me its prayer time see you soon

OpenStudy (anonymous):

you are welcome

OpenStudy (anonymous):

K